OpenStudy (anonymous):

I have a trouble in trigonomerty. Can you help me with this problem cos(square) a + sin(square) a = cos 2a

sam (.sam.):

Proving problem?

sam (.sam.):

\[\cos^2(a) + \sin^2(a) = \cos (2a)\]?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

no, typo thats not square, but 4

sam (.sam.):

\[\cos^4(a) + \sin^4(a) = \cos (2a) \]?

OpenStudy (anonymous):

yeah, that's right can you solve that?